# Quantum metrology of low frequency electromagnetic modes with frequency upconverters

### v2

@@ I Introduction
The field of Circuit Quantum Electrodynamics (Circuit QED) has made impressive strides in harnessing the quantum-mechanical properties of superconducting circuits operating in the microwave frequency regime (typically several GHz) Blais et al. (2007). The techniques of Circuit QED have advanced to the point that detecting Johnson et al. (2010); Chen et al. (2011) and coherently manipulating Nakamura et al. (1999) a single microwave quantum are routine operations, and individual control over arrays of dozens of interacting quantum circuits is possible Arute et al. (2019). Much of this progress has been driven by the desire to build a universal quantum computer capable of performing calculations that would be impractical on any classical computer.
@@
The techniques of Circuit QED do not extend directly to lower frequencies, however. Recently, there has been growing interest in adapting quantum metrology techniques to lower frequency electromagnetic modes, typically at frequencies between dc and the Very High Frequency (VHF) band below 300 MHz. Quantum metrology of low-frequency modes could offer a sensitivity advantage over classical sensors, enabling experiments including dark-matter searches Brouwer et al. (2022), low-frequency nuclear spin metrology Sleator et al. (1985), some astronomical measurements Zmuidzinas (2003), and low-frequency magnetometry, outperforming a dc SQUID in certain applications.
@@
One approach for quantum RF metrology that has been developed at these frequencies is to use a qubit to cool the MHz resonator to its ground state and stabilize a Fock state with a small number of photons Gely et al. (2019). While this approach can be used in principle to measure individual signal photons entering the resonator, it does not provide a way to discriminate incoming signal photons from background thermal photons, which limits its usefulness for certain measurements.
@@
For example, searches for axion or axion-like dark matter at mass below 1 µeV must detect or rule out yoctowatt-scale or smaller electromagnetic signals over many decades in frequency, spanning from $\sim$100Hz to $\sim$300MHz Chaudhuri et al. (2015); Phipps et al. (2020); Ouellet et al. (2019); Crisosto et al. (2020); Garcon et al. (2018); Nguyen et al. (2019); Brouwer et al. (2022). These signals can be used to excite an electromagnetic resonator. If the photon number state of the resonator is then measured, the signal-to-noise ratio is limited by the random entry or departure of background thermal photons from the resonator, since photon-counting techniques lack the frequency resolution to distinguish thermal and signal photons. An alternative approach that applies quantum techniques to measure the dark-matter-induced voltage, rather than the photon number, allows the signal frequency to be determined. Frequency information is important as these circuits carry useful information significantly detuned from their resonant frequency. Far from the resonant frequency, thermal fluctuations are suppressed to below the level of a single photon per second per Hz of bandwidth Chaudhuri et al. (2018, 2019). This off-resonant signal information can be accessed using continuous-variables readout techniques operating beyond the Standard Quantum Limit (SQL). In this case, improving the readout performance does not substantially improve the signal to noise ratio (SNR) on resonance (which is limited by thermal fluctuations), but it allows constant SNR to be maintained over a much broader bandwidth, dramatically increasing the axion search rate.
@@
This upconversion paradigm allows the RQU to take advantage of several mature microwave Circuit QED technologies, including high coherence microwave resonators Reagor et al. (2016), Josephson Parametric Amplifiers (JPAs) Roy and Devoret (2016), and microwave squeezers Castellanos-Beltran et al. (2008), while extending the frequency range of quantum measurement techniques to lower frequencies. They also enable the use of quantum-limit evading techniques equivalent to those used in cavity optomechanics, including back-action evading (BAE) measurementsClerk et al. (2008). After analyzing BAE using the RQU, in §$\S$§VI we experimentally demonstrate a phase-sensitive gain of $46.9$ dB, which is a significant step towards the realization of full BAE.
@@ II.1 Upconverter Hamiltonian
Cavity optomechanics treats two bosonic modes at different frequencies: an electromagnetic mode at $\omega_{a}$ and a mechanical mode at $\omega_{b}$, with $\omega_{a}\gg\omega_{b}$ Aspelmeyer et al. (2014). The position operator of the mechanical mode represents the position of a movable mirror that forms one end of the optical cavity. The modes are quantized with ladder operators $\hat{a},\hat{a}^{\dagger}$ and $\hat{b}$, $\hat{b}^{\dagger}$, respectively. The uncoupled Hamiltonian is:
@@
Figure 1: A circuit model for an RQU, which inductively couples a dc-VHF signal source (shown here as a resonator formed by $C_{b}$ and $L_{b}$) to a tunable Josephson inductance $L_{J}(\hat{\Phi})$. The tunable inductor is made up of a superconducting interferometer with one or more Josephson junctions (JJs) and one or more loops. The flux $\Phi$ threading the inductor $L_{b}$ associated with the low-frequency mode also couples through a designable mutual inductance to each of the loops in the JJ interferometer. Thus, $\hat{\Phi}$ changes the inductance $L_{J}$ presented by the JJ interferometer to the microwave mode, and modifies the resonance frequency of the microwave resonator formed by the interferometer and linear reactances modeled by circuit elements $C_{a}$ and $L_{a}$. A coupling capacitance $C_{c}$, microwave transmission lines and a circulator allow the state of the microwave resonator to be driven and detected via traveling wave modes $\hat{a}_{in}$ and $\hat{a}_{out},$ respectively. The output mode contains information in sidebands, as shown schematically in the frequency domain. Low noise amplification by a cryogenic microwave amplifier allows for efficient detection of the output state $\hat{a}_{out}$.
An optomechanical-style coupling can be realized in microwave superconducting resonant circuits by including a Josephson interferometer whose inductance $L_{J}(\hat{\Phi})$ depends on the flux $\hat{\Phi}$ in the low-frequency mode. The flux causes the microwave resonance frequency to vary, just as position shifts of the moving mirror cause the optical frequency to vary in the optomechanical setup. This optomechanical-style coupling has been realized in the microwave SQUID multiplexer Mates et al. (2008) although not optimized to approach quantum limits. The dispersive nanoSQUID magnetometer Levenson-Falk et al. (2013) uses a similar frequency-tunable microwave resonator, but does not use a resonant low-frequency circuit on its input, removing the effects of quantum backaction and limiting the quantum protocols which could be employed. There are also other devices in which a resonator is tuned with a Josephson junction array, including the Asymmetrically Threaded SQUID (ATS) Lescanne et al. (2020). However, in the ATS, the lower frequency signal and higher frequency signal are both coupled as a current drive to the interferometer, and a flux input is used to bias and pump the interferometer for degenerate four-wave mixing. In the RQU, which uses optimechanical-style coupling, the low-frequency signal is applied as a flux to the Josephson interferometer and three-wave mixing is realized. Figure 1 shows a circuit model of such an upconverter with the associated ladder operators.
@@
In order to treat the interaction between the modes, we include the perturbation of $L_{J}$ due to the sensor flux. For small fluxes satisfying $|\hat{\Phi}|\ll\Phi_{0}$, we Taylor expand the microwave frequency to first order in the sensor flux, to calculate the shift of the microwave resonance frequency due to flux in the sensor:
@@
$\hat{H}_{int}^{RQU}=-\frac{\hbar g_{0}}{\Phi_{ZPF}}\hat{a}^{\dagger}\hat{a}\hat{\Phi}=-\hbar g_{0}\hat{a}^{\dagger}\hat{a}(\hat{b}^{\dagger}+\hat{b}).$ (9)
@@
$g_{0}\equiv\frac{d\omega_{a}}{d\Phi}\Phi_{ZPF}.$ (10)
@@ II.2 Input-Output Model
$\hat{H}_{\rm{tot}}=\hat{H}_{0}+\hat{H}_{\rm{int}}+\hat{H}_{\kappa}+\hat{H}_{\gamma}+\hat{H}_{\rm{drive}},$ (11)
where $\hat{H}_{0}+\hat{H}_{\rm{int}}$ describes the dynamics of the isolated RQU system (microwave resonator plus low-frequency resonator, and their interaction), as described in equations 1 and 7. $\hat{H}_{\kappa}$ captures the effects of loss in the microwave resonator, which is dominated by loss to the strongly coupled readout port. $\hat{H}_{\gamma}$ describes loss to internal dissipation in the low-frequency resonator. Finally, $\hat{H}_{\rm{drive}}$ accounts for the energy supplied by the external drive tones which probe the RQU state.
The traveling-wave modes used in this section are shown in figure 1: the microwave resonator is coupled to an “input” mode $\hat{a}_{in}$ and an “output” mode $\hat{a}_{out}$ which are used to drive and detect the state of the microwave resonator. The field circulating within the microwave resonator, $\hat{a}$, is referred to as the intra-cavity field. A circulator prevents leftward-propagating modes from interacting with the RQU, so we ignore them. The internal dissipation in the low-frequency resonator is modeled as arising from a semi-infinite transmission line of characteristic impedance $R_{b}$. The incident and reflected modes on this transmission line are $\hat{b}_{in}$ and $\hat{b}_{out}$, respectively.
The noise fluctuations in these input and output modes can be analyzed using standard input-output theory Walls and Milburn (2008), following a treatment similar to Section IIIB of Aspelmeyer et al. (2014). The Heisenburg-Langevin equations of motion for the system are:
$\dot{\hat{a}}=\frac{i}{\hbar}\left[\hat{H}_{0}+\hat{H}_{\rm{int}},\hat{a}\right]-\frac{\kappa}{2}\hat{a}-\sqrt{\kappa}\hat{a}_{in},$ (12)
$\dot{\hat{b}}=\frac{i}{\hbar}\left[\hat{H}_{0}+\hat{H}_{\rm{int}},\hat{b}\right]-\frac{\gamma}{2}\hat{b}-\sqrt{\gamma}\hat{b}_{in},$ (13)
where $\kappa$ is the decay rate of the circulating power in the microwave resonator, and $\gamma$ is the decay rate in the low-frequency resonator. Dots indicate time derivatives. $\hat{H}_{int}$ describes three-wave mixing, so equations 12 and 13 are nonlinear. Our present analysis will focus on the regime where we can linearize $\hat{H}_{int}$, although experiments in the highly nonlinear regime (when the single-photon interaction rate $g_{0}$ exceeds the microwave loss rate $\kappa$) are also interesting. To linearize, we write the microwave ladder operators as a sum of a classical amplitude and small quantum fluctuations:
@@
$\hat{a}_{in}=\bar{a}_{in}+\delta\hat{a}_{in},$ (15)
$\hat{a}_{out}=\bar{a}_{out}+\delta\hat{a}_{out}.$ (16)
Here, $|\bar{a}|^{2}$ represents the average photon number circulating in the microwave resonator due to the drive. The drive has an amplitude $|\bar{a}_{in}|^{2}$ (in units of photons/second). Likewise, the average photon flux propagating towards the microwave amplifier is given by $|\bar{a}_{out}|^{2}$. The boundary condition relating the output, input, and intra-cavity fields is:
$(\hat{a}_{out}-\hat{a}_{in})=\sqrt{\kappa}\hat{a}$ (17)
We operate the upconverter in the regime of strong microwave drives, such that $\bar{a}\gg 1$. Inserting expression 14 into the interaction Hamiltonian 7 yields:
$\hat{H}_{int}=-\hbar g_{0}\left(|\bar{a}|^{2}+\bar{a}^{*}\delta\hat{a}+\bar{a}\delta\hat{a}^{\dagger}+\delta\hat{a}^{\dagger}\delta\hat{a}\right)(\hat{b}^{\dagger}+\hat{b}).$ (18)
@@
$\hat{H}_{int}\approx-\hbar g_{0}\left(\bar{a}^{*}\delta\hat{a}+\bar{a}\delta\hat{a}^{\dagger}\right)(\hat{b}^{\dagger}+\hat{b}).$ (19)
Since $\bar{a}$ is a classical value rather than an operator, this interaction now describes an effective two-wave interaction with a tunable strength set by $g_{0}|\bar{a}|$, rather than the three-wave interaction described by equation 7. Hence, we refer to it as the linearized interaction.
@@
$\hat{H}_{\rm{drive}}=-i\hbar\sqrt{\kappa}(\bar{a}_{in}(t)\hat{a}^{\dagger}+\bar{a}_{in}^{*}(t)\hat{a}),$ (20)
where $\bar{a}_{in}(t)$ represents the (classical) amplitude of the coherent voltage drive applied to the microwave resonator via the input transmission line. We apply the unitary transformation:
$\hat{H}_{rot}=\hat{U}\hat{H}_{lab}\hat{U}^{\dagger}-i\hat{U}\dot{\hat{U}}^{\dagger},$ (21)
to move to a frame rotating at the drive frequency. $\hat{U}=\rm{exp}(i\omega_{d}\hat{a}^{\dagger}\hat{a}t)$, where $\omega_{d}$ is the angular velocity of the rotating frame. In this frame, the Heisenberg-Langevin equation of motion for $\hat{a}$ reads:
$\dot{\hat{a}}=\frac{i}{\hbar}\left[-\hbar\left(\Delta+\hbar g_{0}(\hat{b}^{\dagger}+\hat{b})\right)\hat{a}^{\dagger}\hat{a},\hat{a}\right]-\frac{\kappa}{2}\hat{a}-\sqrt{\kappa}\hat{a}_{in},$ (22)
@@
$\bar{a}=\frac{\sqrt{\kappa}\bar{a}_{in}}{i\Delta-\frac{\kappa}{2}}.$ (23)
@@
$-i\omega\delta\hat{a}[\omega]=(i\Delta-\kappa/2)\delta\hat{a}[\omega]-\\ ig_{0}\bar{a}(\hat{b}^{\dagger}[\omega]+\hat{b}[\omega])-\sqrt{\kappa}\delta\hat{a}_{in}[\omega]\$ (24)
@@
$-i\omega\hat{b}[\omega]=(-i\omega_{b}-\gamma/2)\hat{b}[\omega]-\\ ig_{0}(\bar{a}^{*}\delta\hat{a}[\omega]+\bar{a}\delta\hat{a}^{\dagger}[\omega])-\sqrt{\gamma}\hat{b}_{in}[\omega]$ (25)
@@
Equations 24 and 25 fully describe the dynamics of the RQU in the linearized regime, and can be used to calculate the behavior of the coupled modes in a variety of regimes. The quantum metrology techniques in the following sections arise from special cases of these linearized dynamics where the detuning takes on the specific values $\Delta=0$ or a superposition of drive tones at $\Delta=\pm\omega_{b}$.

### v3

@@ I Introduction
The field of Circuit Quantum Electrodynamics (Circuit QED) has made impressive strides in harnessing the quantum-mechanical properties of superconducting circuits operating in the microwave frequency regime (typically several GHz) Blais et al., (2007). The techniques of Circuit QED have advanced to the point that detecting Johnson et al., (2010); Chen et al., (2011) and coherently manipulating Nakamura et al., (1999) a single microwave quantum are routine operations, and individual control over arrays of dozens of interacting quantum circuits is possible Arute et al., (2019). Much of this progress has been driven by the desire to build a universal quantum computer capable of performing calculations that would be impractical on any classical computer.
@@
The techniques of Circuit QED do not extend directly to lower frequencies, however. Recently, there has been growing interest in adapting quantum metrology techniques to lower frequency electromagnetic modes, typically at frequencies between dc and the Very High Frequency (VHF) band below 300 MHz. Quantum metrology of low-frequency modes could offer a sensitivity advantage over classical sensors, enabling experiments including dark-matter searches Brouwer et al., (2022), low-frequency nuclear spin metrology Sleator et al., (1985), some astronomical measurements Zmuidzinas, (2003), and low-frequency magnetometry, outperforming a dc SQUID in certain applications.
@@
One approach for quantum RF metrology that has been developed at these frequencies is to use a qubit to cool the MHz resonator to its ground state and stabilize a Fock state with a small number of photons Gely et al., (2019). While this approach can be used in principle to measure individual signal photons entering the resonator, it does not provide a way to discriminate incoming signal photons from background thermal photons, which limits its usefulness for certain measurements.
@@
For example, searches for axion or axion-like dark matter at mass below 1 µeV must detect or rule out yoctowatt-scale or smaller electromagnetic signals over many decades in frequency, spanning from $\sim$100Hz to $\sim$300MHz Chaudhuri et al., (2015); Phipps et al., (2020); Ouellet et al., (2019); Crisosto et al., (2020); Garcon et al., (2018); Nguyen et al., (2019); Brouwer et al., (2022). These signals can be used to excite an electromagnetic resonator. If the photon number state of the resonator is then measured, the signal-to-noise ratio is limited by the random entry or departure of background thermal photons from the resonator, since photon-counting techniques lack the frequency resolution to distinguish thermal and signal photons. An alternative approach that applies quantum techniques to measure the dark-matter-induced voltage, rather than the photon number, allows the signal frequency to be determined. Frequency information is important as these circuits carry useful information significantly detuned from their resonant frequency. Far from the resonant frequency, thermal fluctuations are suppressed to below the level of a single photon per second per Hz of bandwidth Chaudhuri et al., (2018, 2019). This off-resonant signal information can be accessed using continuous-variables readout techniques operating beyond the Standard Quantum Limit (SQL). In this case, improving the readout performance does not substantially improve the signal to noise ratio (SNR) on resonance (which is limited by thermal fluctuations), but it allows constant SNR to be maintained over a much broader bandwidth, dramatically increasing the axion search rate.
@@
This upconversion paradigm allows the RQU to take advantage of several mature microwave Circuit QED technologies, including high coherence microwave resonators Reagor et al., (2016), Josephson Parametric Amplifiers (JPAs) Roy and Devoret, (2016), and microwave squeezers Castellanos-Beltran et al., (2008), while extending the frequency range of quantum measurement techniques to lower frequencies. They also enable the use of quantum-limit evading techniques equivalent to those used in cavity optomechanics, including back-action evading (BAE) measurementsClerk et al., (2008). After analyzing BAE using the RQU, in §$\S$§VI we experimentally demonstrate a phase-sensitive gain of $46.9$ dB, which is a significant step towards the realization of full BAE.
@@ II.1 Upconverter Hamiltonian
Cavity optomechanics treats two bosonic modes at different frequencies: an electromagnetic mode at $\omega_{a}$ and a mechanical mode at $\omega_{b}$, with $\omega_{a}\gg\omega_{b}$ Aspelmeyer et al., (2014). The position operator of the mechanical mode represents the position of a movable mirror that forms one end of the optical cavity. The modes are quantized with ladder operators $\hat{a},\hat{a}^{\dagger}$ and $\hat{b}$, $\hat{b}^{\dagger}$, respectively. The uncoupled Hamiltonian is:
@@
Figure 1: A circuit model for an RQU, which inductively couples a dc-VHF signal source (shown here as a resonator formed by $C_{b}$ and $L_{b}$) to a tunable Josephson inductance $L_{J}(\hat{\Phi})$. The tunable inductor is made up of a superconducting interferometer with one or more Josephson junctions (JJs) and one or more loops. The flux $\Phi$ threading the inductor $L_{b}$ associated with the low-frequency mode also couples through a designable mutual inductance to each of the loops in the JJ interferometer. Thus, $\hat{\Phi}$ changes the inductance $L_{J}$ presented by the JJ interferometer to the microwave mode, and modifies the resonance frequency of the microwave resonator formed by the interferometer and linear reactances modeled by circuit elements $C_{a}$ and $L_{a}$. A coupling capacitance $C_{c}$, microwave transmission lines and a circulator allow the state of the microwave resonator to be driven and detected via traveling wave modes $\hat{a}_{\rm{in}}$ and $\hat{a}_{\rm{out}},$ respectively. The output mode contains information in sidebands, as shown schematically in the frequency domain. Low noise amplification by a cryogenic microwave amplifier allows for efficient detection of the output state $\hat{a}_{\rm{out}}$.
An optomechanical-style coupling can be realized in microwave superconducting resonant circuits by including a Josephson interferometer whose inductance $L_{J}(\hat{\Phi})$ depends on the flux $\hat{\Phi}$ in the low-frequency mode. The flux causes the microwave resonance frequency to vary, just as position shifts of the moving mirror cause the optical frequency to vary in the optomechanical setup. This optomechanical-style coupling has been realized in the microwave SQUID multiplexer Mates et al., (2008) although not optimized to approach quantum limits. The dispersive nanoSQUID magnetometer Levenson-Falk et al., (2013) uses a similar frequency-tunable microwave resonator, but does not use a resonant low-frequency circuit on its input, removing the effects of quantum backaction and limiting the quantum protocols which could be employed. There are also other devices in which a resonator is tuned with a Josephson junction array, including the Asymmetrically Threaded SQUID (ATS) Lescanne et al., (2020). However, in the ATS, the lower frequency signal and higher frequency signal are both coupled as a current drive to the interferometer, and a flux input is used to bias and pump the interferometer for degenerate four-wave mixing. In the RQU, which uses optimechanical-style coupling, the low-frequency signal is applied as a flux to the Josephson interferometer and three-wave mixing is realized. Figure 1 shows a circuit model of such an upconverter with the associated ladder operators.
@@
In order to treat the interaction between the modes, we include the perturbation of $L_{J}$ due to the sensor flux. We analyze the behavior of the RQU in response to small, time-varying flux signals satisfying $|\langle\hat{\Phi}(t)\rangle|\ll\Phi_{0}$, where $\Phi_{0}=h/2e$ is the magnetic flux quantum that sets the periodicity of the interferometer’s response to external magnetic flux. In this regime, we Taylor expand the microwave frequency to first order in the sensor flux, to calculate the shift of the microwave resonance frequency due to flux in the sensor:
@@
$\hat{H}_{\mathrm{int}}^{\mathrm{RQU}}=-\frac{\hbar g_{0}}{\Phi_{\rm{ZPF}}}\hat{a}^{\dagger}\hat{a}\hat{\Phi}=-\hbar g_{0}\hat{a}^{\dagger}\hat{a}(\hat{b}^{\dagger}+\hat{b}).$ (9)
@@
$g_{0}\equiv\frac{d\omega_{a}}{d\Phi}\Phi_{\rm{ZPF}}.$ (10)
@@ II.2 Input-Output Model
$\hat{H}_{\rm{tot}}=\hat{H}_{0}+\hat{H}_{\rm{int}}^{\rm{RQU}}+\hat{H}_{\kappa}+\hat{H}_{\gamma}+\hat{H}_{\rm{drive}},$ (11)
where $\hat{H}_{0}+\hat{H}_{\rm{int}}^{\rm{RQU}}$ describes the dynamics of the isolated RQU system (microwave resonator plus low-frequency resonator, and their interaction), as described in equations 1 and 7. $\hat{H}_{\kappa}$ captures the effects of loss in the microwave resonator, which is dominated by loss to the strongly coupled readout port. $\hat{H}_{\gamma}$ describes loss to internal dissipation in the low-frequency resonator. Finally, $\hat{H}_{\rm{drive}}$ accounts for the energy supplied by the external drive tones which probe the RQU state.
The traveling-wave modes used in this section are shown in figure 1: the microwave resonator is coupled to an “input” mode $\hat{a}_{\rm{in}}$ and an “output” mode $\hat{a}_{\rm{out}}$ which are used to drive and detect the state of the microwave resonator. The field circulating within the microwave resonator, $\hat{a}$, is referred to as the intra-cavity field. A circulator prevents leftward-propagating modes from interacting with the RQU, so we ignore them. The internal dissipation in the low-frequency resonator is modeled as arising from a semi-infinite transmission line of characteristic impedance $R_{b}$. The incident and reflected modes on this transmission line are $\hat{b}_{\rm{in}}$ and $\hat{b}_{\rm{out}}$, respectively.
The noise fluctuations in these input and output modes can be analyzed using standard input-output theory Aspelmeyer et al., (2014); Walls and Milburn, (2008), treating $\hat{H}_{\kappa}$ and $\hat{H}_{\gamma}$ as the microwave and low-frequency bath Hamiltonians, respectively. After making the Markovian approximation neglecting memory effects in the bath modes, the Heisenburg-Langevin equations of motion for the system are:
$\dot{\hat{a}}=\frac{i}{\hbar}\left[\hat{H}_{0}+\hat{H}_{\rm{int}}^{\rm{RQU}},\hat{a}\right]-\frac{\kappa}{2}\hat{a}-\sqrt{\kappa}\hat{a}_{\rm{in}},$ (12)
$\dot{\hat{b}}=\frac{i}{\hbar}\left[\hat{H}_{0}+\hat{H}_{\rm{int}}^{\rm{RQU}},\hat{b}\right]-\frac{\gamma}{2}\hat{b}-\sqrt{\gamma}\hat{b}_{\rm{in}},$ (13)
where $\kappa$ is the decay rate of the circulating power in the microwave resonator, and $\gamma$ is the decay rate in the low-frequency resonator. Dots indicate time derivatives. $\hat{H}_{\rm{int}}^{\rm{RQU}}$ describes three-wave mixing, so equations 12 and 13 are nonlinear. Our present analysis will focus on the regime where we can linearize $\hat{H}_{\mathrm{int}}$, although experiments in the highly nonlinear regime (when the single-photon interaction rate $g_{0}$ exceeds the microwave loss rate $\kappa$) are also interesting. To linearize, we write the microwave ladder operators as a sum of a classical amplitude and small quantum fluctuations:
@@
$\hat{a}_{\rm{in}}=\bar{a}_{\rm{in}}+\delta\hat{a}_{\rm{in}},$ (15)
$\hat{a}_{\rm{out}}=\bar{a}_{\rm{out}}+\delta\hat{a}_{\rm{out}}.$ (16)
Here, $|\bar{a}|^{2}$ represents the average photon number circulating in the microwave resonator due to the drive. The drive has an amplitude $|\bar{a}_{\rm{in}}|^{2}$ (in units of photons/second). Likewise, the average photon flux propagating towards the microwave amplifier is given by $|\bar{a}_{\rm{out}}|^{2}$. The boundary condition relating the output, input, and intra-cavity fields is:
$(\hat{a}_{\rm{out}}-\hat{a}_{\rm{in}})=\sqrt{\kappa}\hat{a}$ (17)
We operate the upconverter in the regime of strong microwave drives, such that $|\bar{a}|\gg 1$. Inserting expression 14 into the interaction Hamiltonian 9 yields:
$\hat{H}_{\mathrm{int}}^{\rm{RQU}}=-\hbar g_{0}\left(|\bar{a}|^{2}+\bar{a}^{*}\delta\hat{a}+\bar{a}\delta\hat{a}^{\dagger}+\delta\hat{a}^{\dagger}\delta\hat{a}\right)(\hat{b}^{\dagger}+\hat{b}).$ (18)
@@
$\hat{H}_{\rm{int}}^{\rm{RQU}}\approx-\hbar g_{0}\left(\bar{a}^{*}\delta\hat{a}+\bar{a}\delta\hat{a}^{\dagger}\right)(\hat{b}^{\dagger}+\hat{b}).$ (19)
Since $\bar{a}$ is a classical value rather than an operator, this interaction now describes an effective two-wave interaction with a tunable strength set by $g_{0}|\bar{a}|$, rather than the three-wave interaction described by equation 7. This approximation holds for strong microwave drivesHence, we refer to it as the linearized interaction.
@@
$\hat{H}_{\rm{drive}}=-i\hbar\sqrt{\kappa}(\bar{a}_{\rm{in}}(t)\hat{a}^{\dagger}+\bar{a}_{\rm{in}}^{*}(t)\hat{a}),$ (20)
where $\bar{a}_{\rm{in}}(t)$ represents the (classical) amplitude of the coherent voltage drive applied to the microwave resonator via the input transmission line. We apply the unitary transformation:
$\hat{H}_{\rm{rot}}=\hat{U}\hat{H}_{\rm{lab}}\hat{U}^{\dagger}-i\hat{U}\dot{\hat{U}}^{\dagger},$ (21)
to move to a frame rotating at the drive frequency. $\hat{U}=e^{(i\omega_{d}\hat{a}^{\dagger}\hat{a}t)}$, where $\omega_{d}$ is the angular velocity of the rotating frame. We use $\hat{H}_{\mathrm{lab}}=\hat{H}_{0}+\hat{H}_{\mathrm{int}}^{\mathrm{RQU}}$ as the lab-frame Hamiltonian. In this frame, the Heisenberg-Langevin equation of motion for $\hat{a}$ reads:
$\dot{\hat{a}}=\frac{i}{\hbar}\left[-\hbar\left(\Delta+g_{0}(\hat{b}^{\dagger}+\hat{b})\right)\hat{a}^{\dagger}\hat{a},\hat{a}\right]-\frac{\kappa}{2}\hat{a}-\sqrt{\kappa}\hat{a}_{\rm{in}},$ (22)
@@
$\bar{a}=\frac{\sqrt{\kappa}\bar{a}_{\rm{in}}}{i\Delta-\frac{\kappa}{2}}.$ (23)
@@
$-i\omega\delta\hat{a}[\omega]=(i\Delta-\kappa/2)\delta\hat{a}[\omega]-\\ ig_{0}\bar{a}(\hat{b}^{\dagger}[\omega]+\hat{b}[\omega])-\sqrt{\kappa}\delta\hat{a}_{\rm{in}}[\omega]\$ (24)
@@
$-i\omega\hat{b}[\omega]=(-i\omega_{b}-\gamma/2)\hat{b}[\omega]-\\ ig_{0}(\bar{a}^{*}\delta\hat{a}[\omega]+\bar{a}\delta\hat{a}^{\dagger}[\omega])-\sqrt{\gamma}\hat{b}_{\rm{in}}[\omega]$ (25)
@@
Equations 24 and 25 fully describe the dynamics of the RQU in the linearized regime, and can be used to calculate the behavior of the coupled modes in a variety of regimes. The approximations used in this analysis hold in the regime of strong microwave drives ($|\bar{a}|\gg 1$) and moderate interaction strength ($g_{0}<\kappa$), and they simplify the dynamics of the coupled system of resonators to an effective two-wave interaction with tunable strength. As discussed in the next section, this description allows the RQU system to be analyzed as a linear, op-amp mode amplifier. The quantum metrology technique in §V arises from a special cases of these linearized dynamics where microwave drive consists of a superposition of drive tones at $\Delta=\pm\omega_{b}$.